# Rethinking Spectral Graph Neural Networks with Spatially Adaptive Filtering

### 17 Jan 2024 v1

@@ 6 Experiments
Full-supervised Node Classification. To bolster our evaluation for the full-supervised node classification task, we augment the baseline models (previously used in semi-supervised node classification) with seven cutting-edge spatial GNNs: GCNII, FAGCN, GEN, PDE-GCN, NodeFormer, and GloGNN++. For all datasets, we randomly divide them into 60%/20%/20% for training/validation/testing by following (He et al., 2021; 2022). Table 2 summarizes the mean classification accuracies. Notably, provided with more training samples, our proposed methods beat the leading spectral and spatial GNNs, except on Squi. dataset where they achieve comparable results to Specformer. This exceptional performance stems from its non-local aggregation, with signed edge weights that globally address both node similarity and dissimilarity. This allows our models to effectively mitigate long-standing issues inherent in GNNs, such as long-range dependencies and graph heterophily. In addition to the eight common datasets for node classification, we also incorporated experimental results from five recently proposed benchmarks tailored for graph heterophily. Due to space limit, the detailed results of these additional benchmarks are presented in Appendix F.2.
(a) Ablation study of SAF on six datasets.
(b) Squirrel
@@
Figure 3: (a) Comparison between SAF and its ablated variant in full-supervised node classification accuracy. (b-c) Changing trends of the averaged attention values with respect to training epochs.
Ablation Study. Despite the critical contribution of non-local aggregation to SAF, spectral filtering remains vital for discriminative node representation learning. Specifically, the quality of the adapted new graph fundamentally hinges on the graph spectral filters’ training, as underscored by their theoretical interaction in Eq. (6). To provide empirical evidences, we conduct an ablation study on SAF by removing spectral filtering (i.e., $\boldsymbol{\kappa_{f}}=\mathbf{0},\boldsymbol{\kappa_{a}}=\mathbf{1}$ in Eq. (7)). From Figure 2(a), the absence of spectral filtering results in a noticeable accuracy drop, confirming the importance of spectral filtering in SAF. For a comprehensive ablation analysis of other SAF modules, refer to Appendix F.3.
Analysis of Attention Trends. We analyze the changing trends of the pair-wise attention weights during training SAF on Squirrel and Cora datasets. From Figure 3, the average weights for filtering and aggregation start similarly but diverge throughout training, showing different trends in heterophilic and homophilic graphs. On the heterophilic graph Squirrel, both weights converge to similar values, demonstrating their mutual importance in modeling complex connectivity. Conversely, $\boldsymbol{\kappa}_{f}$ becomes dominant on the homophilic graph Cora due to the sufficiency of node proximity information for label prediction, thereby diminishing the relevance of $\boldsymbol{\kappa}_{a}$ and non-local aggregation.
7 Conclusion and Future Work
This paper introduces a fresh spatial perspective on spectral GNNs, shedding light on their interpretability. We reveal that spectral GNNs fundamentally leads the original graph to an adapted new one, which exhibits non-locality and accommodates signed edge weights to reflect label consistency among nodes. This insight leads to our proposed Spatially Adaptive Filtering (SAF) framework, enhancing spectral GNNs for more effective and versatile graph representation learning.
While SAF adeptly captures long-range dependencies and addresses graph heterophily, it has limitations. Specifically, its non-negative constraints on graph filters might limit filter expressiveness, indicating room for theoretical refinement. Furthermore, although this study focuses on a node-level investigation, it raises intriguing questions about the implications of spectral GNNs at the graph-level in the spatial domain. Future work could expand this examination by exploring the interplay between spatial and spectral domains from a more comprehensive graph-level viewpoint.
Appendix A Overall Pipeline of Spatially Adaptive Filtering Framework
Figure 4: Illustration of the proposed SAF framework, where varying node colors on the input graph represent different node labels.
Appendix B Proofs
B.1 Proof of Proposition 1
@@
We begin with the assertion that the eigenvalues of $\mathbf{I}-g_{\psi}(\hat{\mathbf{L}})$ are positive and strictly less than 1, which fulfills the necessary condition for the Neumann series expansion stated in Lemma 1. As such, we can deduce $g_{\psi}(\hat{\mathbf{L}})^{-1}=(\mathbf{I}-(\mathbf{I}-g_{\psi}(\hat{\mathbf{L}})))^{-1}=\sum_{t=0}^{\infty}(\mathbf{I}-g_{\psi}(\hat{\mathbf{L}}))^{t}$. Owning to the prevalent polynomial approximation, we are eligible to express $g_{\psi}(\hat{\mathbf{L}})$ w.r.t. adjacency matrix $\hat{\mathbf{A}}$, i.e., $g_{\psi}(\hat{\mathbf{L}})=g_{\psi}(\mathbf{I}-\hat{\mathbf{A}})=\sum^{K}_{k=0}\pi_{k}\hat{\mathbf{A}}^{k}$ where $\pi_{k}$ refers to the new coefficients made of up $\{\psi_{m}\}^{K}_{m=0}$. Substituting this polynomial representation into our Neumann expansion, we obtain $g_{\psi}(\hat{\mathbf{L}})^{-1}=\sum_{n=0}^{\infty}(\mathbf{I}-\sum^{K}_{k=0}\pi_{k}\hat{\mathbf{A}}^{k})^{t}$. Now, revisiting $\hat{\mathbf{A}}^{\text{new}}$ formulated in Eq. (6), we have $\hat{\mathbf{A}}^{\text{new}}=\mathbf{I}-\frac{\alpha}{1-\alpha}(g_{\psi}(\hat{\mathbf{L}})^{-1}-\mathbf{I})=\mathbf{I}-\frac{\alpha}{1-\alpha}\sum_{t=1}^{\infty}(\mathbf{I}-\sum^{K}_{k=0}\pi_{k}\hat{\mathbf{A}}^{k})^{t}=\sum_{t=0}^{\infty}\phi_{t}\hat{\mathbf{A}}^{t}$ where $\phi_{t}$ denotes a constant coefficient made up of $\{\pi_{m}\}^{K}_{m=0}$. ∎
B.2 Proof of Proposition 2
@@
Let us commence the proof by contradiction. Let $\mathcal{C}$ denote the condition described in proposition 2. Assume, for the sake of contradiction, that $\mathcal{C}$ is not requisite for the optimal node representations $\mathbf{Z}^{*}$ to be predictive of node labels. Under this assumption, there are node pairs $v_{i},v_{j}\in\mathcal{V}$ such that: (1) if $\mathbf{y}_{i}=\mathbf{y}_{j}$, $\hat{A}^{\text{new}}_{i,j}<0$; (2) if $\mathbf{y}_{i}\neq\mathbf{y}_{j}$, $\hat{A}^{\text{new}}_{i,j}>0$. Without loss of generality, given the non-locality as proved in proposition 1, we exclude cases where $\hat{A}^{\text{new}}_{i,j}=0$ from our consideration. Now, consider the second objective term $\mathrm{tr}(\mathbf{Z}^{T}\gamma_{\theta}(\hat{\mathbf{L}})\mathbf{Z})$ in Eq. (3). Using the relationship $\gamma_{\theta}(\hat{\mathbf{L}})=\mathbf{I}-\hat{\mathbf{A}}^{\text{new}}$, we can expand this term into $\sum_{v_{i},v_{j}\in\mathcal{V}}\hat{A}^{\text{new}}_{i,j}\|\mathbf{Z}_{i}-\mathbf{Z}_{j}\|^{2}_{2}$. Under (1), for same-class nodes $v_{i},v_{j}$ with $\hat{A}^{\text{new}}_{i,j}<0$, minimizing the objective term pulls $\mathbf{Z}_{i}$ and $\mathbf{Z}_{j}$ apart in the representation space. This behavior violates the canonical understanding that nodes from the same class should exhibit similar representations. Under (2), for different-class nodes $v_{i},v_{j}$ with $\hat{A}^{\text{new}}_{i,j}>0$, the optimization encourages $\mathbf{Z}_{i}$ and $\mathbf{Z}_{j}$ to be more similar. This is in direct opposition to the basic classification principle that nodes from different classes should have distinct representations. Given these contradictions stemming from the mathematical implications in optimization, we must reject assumptions (1) and (2), affirming the necessary condition $\mathcal{C}$ for accurate label prediction by $\mathbf{Z}^{*}$. ∎
B.3 Proof of Proposition 3
@@
$p(x)=\sum_{k=0}^{K}\psi_{k}\binom{K}{k}(1-x)^{K-k}x^{k}\leq\psi_{\text{max}}\sum_{k=0}^{K}\binom{K}{k}(1-x)^{K-k}x^{k}=\psi_{\text{max}}(1-x+x)^{K}=\psi_{\text{max}}.$
@@
Appendix C Further Remarks about Related Work
C.1 Polynomial Spectral Graph Neural Networks
Vanilla GCN (Kipf & Welling, 2017). The vanilla GCN truncates Chebyshev polynomials to a simple first-order for efficient graph convolution, which functions as a low-pass filter, i.e., $\mathbf{Z}=w(\mathbf{I}+\mathbf{D}^{-\frac{1}{2}}\mathbf{A}\mathbf{D}^{-\frac{1}{2}})\mathbf{X}$. To ensure numerical stability, it further leverages a renormalization trick to replace $\mathbf{I}+\mathbf{D}^{-\frac{1}{2}}\mathbf{A}\mathbf{D}^{-\frac{1}{2}}$ with $(\mathbf{D}+\mathbf{I})^{-\frac{1}{2}}(\mathbf{A}+\mathbf{I})(\mathbf{D}+\mathbf{I})^{-\frac{1}{2}}$ and derives each graph convolution layer as
$\mathbf{Z}^{(k+1)}=\tilde{\mathbf{D}}^{-\frac{1}{2}}\tilde{\mathbf{A}}\tilde{\mathbf{D}}^{-\frac{1}{2}}\mathbf{Z}^{(k)}\mathbf{W}^{(k)},$
where $\tilde{\mathbf{A}}=\mathbf{A}+\mathbf{I}$, $\tilde{\mathbf{D}}=\mathbf{D}+\mathbf{I}$, $\mathbf{Z}^{(0)}=\mathbf{X}$, and $\mathbf{W}^{(k)}$ denotes learnable weights in the $k$-th layer.
GPR-GNN (Chien et al., 2021). GPR-GNN leverages the Generalized PageRank to approximate spectral graph filters with Monomial polynomial basis. The model architecture is formulated as
$\mathbf{Z}=\sum_{k=0}^{K}\psi_{k}\hat{\tilde{\mathbf{A}}}^{k}f_{\varphi}(\mathbf{X}),$
where $\hat{\tilde{\mathbf{A}}}=\tilde{\mathbf{D}}^{\frac{1}{2}}\tilde{\mathbf{A}}\tilde{\mathbf{D}}^{\frac{1}{2}}$, and $f_{\varphi}(\mathbf{X})$ refers to a linear map with parameters $\varphi$. The approximated graph filter is $g_{\psi}(\tilde{\lambda})=\sum_{k=0}^{K}\psi_{k}\tilde{\lambda}^{k}$ with $\tilde{\lambda}$ being the eigenvalues of $\hat{\tilde{\mathbf{A}}}$.
BernNet (He et al., 2021). BernNet approximates graph filters with Bernstein polynomials to attain non-negative property, yielding the model expression as
$\mathbf{Z}=\sum_{k=0}^{K}\psi_{k}\frac{1}{2^{K}}\binom{K}{k}(2\mathbf{I}-\hat{\mathbf{L}})^{K-k}\hat{\mathbf{L}}^{k}f_{\varphi}(\mathbf{X}).$
The graph filter is defined as $g_{\psi}(\lambda)=\sum_{k=0}^{K}\psi_{k}\frac{1}{2^{K}}\binom{K}{k}(2-\lambda)^{K-k}\lambda^{k}$ and $\lambda$ is the eigenvalue of $\hat{\mathbf{L}}$.
ChebNetII (He et al., 2022). ChebNetII leverages Chebyshev interpolation to approximate graph filter while reducing the Runge phenomenon. This results in the following model structure:
$\mathbf{Z}=\mathbf{Y}=\frac{2}{K+1}\sum_{k=0}^{K}\sum_{j=0}^{K}\gamma_{j}T_{k}\left(x_{j}\right)T_{k}(\hat{\mathbf{L}})f_{\varphi}(\mathbf{X}).$
Here, $T_{k}(x)=2xT_{k-1}(x)-T_{k-2}(x)$ denotes the Chebyshev basis with $T_{0}(x)=1$ and $T_{1}(x)=x$, and $x_{j}=\mathrm{cos}(\frac{(j+1/2)\pi}{K+1})$ refers to the Chebyshev nodes of $T_{K+1}(x)(\cdot)$. The graph filter corresponds to $g_{\psi}(\lambda)=\frac{2}{K+1}\sum_{k=0}^{K}\psi_{k}T_{k}(\lambda)$ with $\psi_{k}=\sum_{j=0}^{K}\gamma_{j}T_{k}\left(x_{j}\right)$.
JacobiConv (Wang & Zhang, 2022). JacobiConv approximates graph filters using orthogonal Jacobi polynomial basis and feature-wise filter learning. It can be expressed as
$\mathbf{Z}_{[:,l]}=\sum_{k=0}^{K}\psi_{k,l}\sum_{s=0}^{k}\beta_{k,s}(2\mathbf{I}-\hat{\mathbf{L}})^{s}(-\hat{\mathbf{L}})^{k-s}\mathbf{X}\mathbf{W}_{[:,l]},$
where $\beta_{k,s}=\frac{(k+a)!(k+b)!}{2^{k}s!(k+a-s)!(b+s)!(k-s)!}$ is a constant coefficient with hyper-parameters $a,b$ and $l$ denotes the $l$-th feature channel. Thus, the feature-wise graph filter is given as $g_{\psi,l}(\lambda)=\sum_{k=0}^{K}\psi_{k,l}\sum_{s=0}^{k}\beta_{k,s}(2-\lambda)^{s}(-\lambda)^{k-s}$.
Despite the proliferation of various filtering strategies, they invariably resort to approximating graph filters with fixed-order polynomials. While this approach does offer computational advantages, the truncated approximation limits the effective propagation range and hinder the ability to capture the long-range dependencies.
C.2 Intertwined Issues of Long-range Dependencies and Graph Heterophily
Long-range dependencies and graph heterophily are intrinsically connected. The necessity to capture long-range dependencies often arises in heterophilic (disassortative) graphs, while the phenomenon of graph heterophily further intensifies the challenges associated with long-range dependencies. In light of this, we aim to conduct a thorough evaluation of our model, comparing it to methods (except for polynomial spectral GNNs) that are designed with singular or dual focus on capturing long-range dependencies and addressing graph heterophily. From the plethora of models extant in the literature, we carefully select a number of representative and advanced methods for comparison. These include three methods for solely capturing long-range dependencies: APPNP (Gasteiger et al., 2019), GCNII (Chen et al., 2020), PDE-GCN (Eliasof et al., 2021), two models focusing on tackling graph heterophily: GEN (Wang et al., 2021), FAGCN (Bo et al., 2021), as well as two works considering both: NodeFormer (Wu et al., 2022), GloGNN++ (Li et al., 2022).
Table 3: Statistics of real-world datasets. Diameter refers to the longest geodesic distance between nodes on the graph. Both $\mathcal{H}$, $\mathcal{H}_{\text{class}}$, and $\mathcal{H}_{\text{adjusted}}$ measure graph homophily ratio from 0 to 1.
Dataset # Nodes # Edges # Features # Classes Diameter $\mathcal{H}$ $\mathcal{H}_{\text{class}}$ $\mathcal{H}_{\text{adjusted}}$
@@ C.2 Intertwined Issues of Long-range Dependencies and Graph Heterophily
Twtich-DE 9,498 153,138 2,514 2 7 0.63 0.14 0.14
C.3 Practical Significance of eigendecomposition

### 08 Feb 2024 v3

@@ 6.2.2. Full-supervised Node Classification
To bolster our evaluation, we expand the previously compared baselines to include six cutting-edge spatial GNN models: GCNII & PDE-GCN for capturing long-range dependency, GEN & MGNN for handling graph heterophily, and NodeFormer & GloGNN++ addressing both. For all datasets, we randomly divide them into 60%/20%/20% for training/validation/testing by following (He et al., 2021, 2022). Table 2 summarizes the mean classification accuracies. Our methods demonstrate superior performance across most datasets, with an exception on Squirrel where they achieve comparable results to Specformer. This notable performance is primarily attributed to our SAF’s effective non-local aggregation, utilizing signed edge weights to model global label relationships. This enables our methods to outperform GNNs that are specifically tailored for long-range dependency and/or graph heterophily. Besides the eight standard datasets for node classification, our study also extends to five new benchmarks (Lim et al., 2021; Platonov et al., 2023b) that focus on graph heterophily. Due to space limit, we present detailed results in Appendix E.3.
Figure 4. Ablation study of SAF framework on six datasets. Refer to Table 8 in the Appendix for comprehensive results.
6.3. Ablation Study
This section aims to validate our designs by comparing SAF with its three ablated variants – SAF w/o Atte., SAF w/o Spec., and SAF w/o Spat. – in full-supervised node classification. Specifically, Atte., Spec., and Spat. respectively refers to: attention mechanism in “Node-wise Prediction Amalgamation”, “Non-negative Spectral Filtering”, and “Non-local Spatial Aggregation”. For SAF w/o Atte., we remove the attention mechanism and equally blend predictions from different domains. SAF w/o Spec. abandons the spectral filtering phase, practically setting $\boldsymbol{\kappa_{f}}=\mathbf{0},\boldsymbol{\kappa_{a}}=\mathbf{1}$ in Eq. (7). As the SAF w/o Spat. configuration is equivalent to BernNet model, the corresponding results are posted directly. From Figure 4, we can draw several insights: 1) The impact of Atte. module on our SAF varies by datasets, e.g., on Chameleon and Squirrel, showing a slight performance reduction upon its removal. This observation aligns with our observation that their optimal attention values are close to an even split, as suggested in Figures 5(a) and 9(a) Conversely, Cora dataset exhibits a notable drop, due to its optimal attention weights being far from even, as depicted in Figure 5(b). 2) Spectral filtering (Spec. module) remains vital for discriminative node representation learning. Specifically, the quality of the adapted new graph fundamentally hinges on the graph spectral filters’ training, as underscored by their theoretical interaction in Eq. (6). Practically, the absence of spectral filtering markedly reduces model accuracy, confirming its importance in SAF. 3) This visualization not only reaffirms the pivotal role of the non-local aggregation (Spat. module), but also underscores its position as the most crucial component in advancing spectral GNNs within the SAF framework.
6.4. Analysis of Attention Trends
We analyze the changing trends of the pair-wise attention weights during training SAF on Squirrel and Cora datasets. From Figure 5, the average weights for filtering and aggregation start similarly but diverge throughout training, showing different trends in heterophilic and homophilic graphs. On the heterophilic graph Squirrel, both weights converge to similar values, demonstrating their mutual importance in modeling complex connectivity. Conversely, $\boldsymbol{\kappa}_{f}$ becomes dominant on the homophilic graph Cora due to the sufficiency of node proximity information for label prediction, thereby diminishing the relevance of $\boldsymbol{\kappa}_{a}$ and non-local aggregation.
6.5. Parameter Study
This section presents the sensitivity analysis of hyper-parameters including $\tau$, $\eta$, $\epsilon$, and $L$. Figure 6 visualizes how varying these parameters within a broad range influences learning performance, showcasing our model’s robust stability over diverse settings. Beyond empirical observation, we also provides deeper insights into parameter understanding and rationalizes the chosen ranges for parameter searching: 1) The scaling parameter $\tau=\frac{\alpha}{1-\alpha}$, crucial in new graph construction in Eq. (6), stems from the trade-off parameter $\alpha\in(0,1)$ within the graph optimization problem in Eq.(3). While theoretically we have $0=\frac{0}{1-0}<\tau<\frac{1}{1-1}=\infty$, practical considerations for extracting structural information suggest a larger penalty on the trace objective term $\mathrm{tr}(\mathbf{Z}^{T}\gamma_{\theta}(\hat{\mathbf{L}})\mathbf{Z})$, i.e., keeping $\alpha<0.5$, thereby limiting $\tau<\frac{0.5}{1-0.5}=1$. This rationale substantiates our selection of $\tau$ within the set $\{0.1,0.2,...,1\}$ as stated in Appendix D.2, aligning with the observed optimal performance in Figures 6(a)-(c). When addressing graphs with noisy structure, we may adjust the upper limit of $\alpha$ to $t\in(0,1)$, setting $\tau$’s maximum possible value to $\frac{t}{1-t}$. For graph benchmarking evaluations in this work, where extracting structural information is important, we practically set $t=0.5$. 2) For the non-local aggregation layer number $L$, a noticeable decline in model performance is observed when $L$ exceeds 10. This is attributed to the non-local nature of our new graph, which facilitates efficient information exchange between nodes. Exceeding a certain number of layers may potentially lead to oversmoothing, where there is an overemphasis on global information, thus degrading model performance. However, choosing the number of layers within a reasonable range generally ensures consistent and impressive model performance, as verified in Figures 6(j)-(l).
(a) Squirrel
(b) Cora
Figure 5. Attention changing trends w.r.t. training epochs. Refer to Figure 9 in the Appendix E.2 for more visualizations.
(a) Chameleon
(b) Texas
@@ 6.5. Parameter Study
(d) Chameleon
(e) Texas
(f) Cora
(g) Chameleon
(h) Texas
(i) Cora
(j) Chameleon
(k) Texas
(l) Cora
Figure 6. Sensitivity analysis for hyper-parameters: $\tau$ (blue), $\eta$ (red), $\epsilon$ (green), and $L$ (orange) from top to bottom rows.
7. Conclusion and Future Work
This paper introduces a fresh spatial perspective on spectral GNNs, shedding light on their interpretability. We reveal that spectral GNNs fundamentally leads the original graph to an adapted new one, which exhibits non-locality and accommodates signed edge weights. This insight leads to our proposed Spatially Adaptive Filtering (SAF) framework, enhancing spectral GNNs for more effective and versatile graph representation learning. While this study focuses on a node-level investigation, it raises intriguing questions about the implications of spectral GNNs at the graph-level in the spatial domain. Future work could expand this examination by exploring the interplay between spatial and spectral domains from a more comprehensive graph-level viewpoint.
Appendix A Proofs
A.1. Proof of Proposition 2
@@
We begin with the assertion that the eigenvalues of $\mathbf{I}-g_{\psi}(\hat{\mathbf{L}})$ are positive and strictly less than 1, which fulfills the necessary condition for the Neumann series expansion stated in Lemma 1. As such, we can deduce $g_{\psi}(\hat{\mathbf{L}})^{-1}=(\mathbf{I}-(\mathbf{I}-g_{\psi}(\hat{\mathbf{L}})))^{-1}=\sum_{t=0}^{\infty}(\mathbf{I}-g_{\psi}(\hat{\mathbf{L}}))^{t}$. Owning to the prevalent polynomial approximation, we are eligible to express $g_{\psi}(\hat{\mathbf{L}})$ w.r.t. adjacency matrix $\hat{\mathbf{A}}$, i.e., $g_{\psi}(\hat{\mathbf{L}})=g_{\psi}(\mathbf{I}-\hat{\mathbf{A}})=\sum^{K}_{k=0}\pi_{k}\hat{\mathbf{A}}^{k}$ where $\pi_{k}$ refers to the new coefficients made of up $\{\psi_{m}\}^{K}_{m=0}$. Substituting this polynomial representation into our Neumann expansion, we obtain $g_{\psi}(\hat{\mathbf{L}})^{-1}=\sum_{n=0}^{\infty}(\mathbf{I}-\sum^{K}_{k=0}\pi_{k}\hat{\mathbf{A}}^{k})^{t}$. Now, revisiting $\hat{\mathbf{A}}^{\text{new}}$ in Eq. (6), we have $\hat{\mathbf{A}}^{\text{new}}=\mathbf{I}-\frac{\alpha}{1-\alpha}(g_{\psi}(\hat{\mathbf{L}})^{-1}-\mathbf{I})=\mathbf{I}-\frac{\alpha}{1-\alpha}\sum_{t=1}^{\infty}(\mathbf{I}-\sum^{K}_{k=0}\pi_{k}\hat{\mathbf{A}}^{k})^{t}=\sum_{t=0}^{\infty}\phi_{t}\hat{\mathbf{A}}^{t}$ where $\phi_{t}$ is a constant coefficient made up of $\{\pi_{m}\}^{K}_{m=0}$. ∎
A.2. Proof of Proposition 3
@@
Let us commence the proof by contradiction. Let $\mathcal{C}$ denote the condition described in proposition 3. Assume, for the sake of contradiction, that $\mathcal{C}$ is not requisite for the optimal node representations $\mathbf{Z}^{*}$ to be predictive of node labels. Under this assumption, there are node pairs $v_{i},v_{j}\in\mathcal{V}$ such that: (1) if $\mathbf{y}_{i}=\mathbf{y}_{j}$, $\hat{A}^{\text{new}}_{i,j}<0$; (2) if $\mathbf{y}_{i}\neq\mathbf{y}_{j}$, $\hat{A}^{\text{new}}_{i,j}>0$. Without loss of generality, given the non-locality as proved in proposition 2, we exclude cases where $\hat{A}^{\text{new}}_{i,j}=0$ from our consideration. Now, consider the second objective term $\mathrm{tr}(\mathbf{Z}^{T}\gamma_{\theta}(\hat{\mathbf{L}})\mathbf{Z})$ in Eq. (3). Using the relationship $\gamma_{\theta}(\hat{\mathbf{L}})=\mathbf{I}-\hat{\mathbf{A}}^{\text{new}}$, we can expand this term into $\sum_{v_{i},v_{j}\in\mathcal{V}}\hat{A}^{\text{new}}_{i,j}\|\mathbf{Z}_{i}-\mathbf{Z}_{j}\|^{2}_{2}$. Under (1), for same-class nodes $v_{i},v_{j}$ with $\hat{A}^{\text{new}}_{i,j}<0$, minimizing the objective term pulls $\mathbf{Z}_{i}$ and $\mathbf{Z}_{j}$ apart in the latent space. This behavior violates the canonical understanding that nodes from the same class should exhibit similar representations. Under (2), for different-class nodes $v_{i},v_{j}$ with $\hat{A}^{\text{new}}_{i,j}>0$, the optimization encourages $\mathbf{Z}_{i}$ and $\mathbf{Z}_{j}$ to be more similar. This is in direct opposition to the basic classification principle that nodes from different classes should have distinct representations. Given these contradictions stemming from the mathematical implications in optimization, we must reject assumptions (1) and (2), affirming the necessary condition $\mathcal{C}$ for accurate label prediction by $\mathbf{Z}^{*}$. ∎
A.3. Proof of Proposition 1
@@
$p(x)$ $=\sum_{k=0}^{K}\psi_{k}\binom{K}{k}(1-x)^{K-k}x^{k}\leq\psi_{\text{max}}\sum_{k=0}^{K}\binom{K}{k}(1-x)^{K-k}x^{k}$
$=\psi_{\text{max}}(1-x+x)^{K}=\psi_{\text{max}}.$
@@
Table 3. Statistics of real-world datasets. $F$ and $C$ denotes the number of features and classes. $\Delta$ represents graph diameter referring to the longest geodesic distance between nodes on the graph. For Penn94, due to multiple subgraphs, we report $\Delta$ of the largest connected component.
[…]
@@
Penn94 41,554 1,362,229 5 2 8 0.47 0.05 0.02
Appendix B Eigendecomposition